# Random Projection Layers for Multidimensional Time Series Forecasting

### v2

@@ A.5 Ablation Study
We have produced correlation-error diagrams employing all three measurements (MAE, RMSE, and MAPE). In Section 5.4, we showcase the correlation-error diagram with MAE in Fig. 8. This section introduces the remaining two correlation-error diagrams, namely, the correlation-error diagram with RMSE in Fig. 11(a), and the correlation-error diagram with MAPE in Fig. 11(b). The conclusions remain consistent. The random projection layer aids the proposed method in achieving greater diversity.
Figure 13: The ablation study result measured using per-step MAE, RMSE, and MAPE. The random projection layer benefits the prediction for later time steps more compared to earlier time steps.
@@
A.6 Long-Term Time Series Forecasting
The effectiveness of the proposed method in spatial-temporal forecasting tasks has been demonstrated. However, as a general time series forecasting method, we sought to assess its performance against other forecasting methods. To this end, we evaluated our method on seven multivariate long-term time series forecasting datasets (i.e., ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity, and Traffic), and compared it with two state-of-the-art methods, TSMixer Chen et al. (2023) and PatchTST Nie et al. (2022), which have shown superior performance over alternatives such as Autoformer Wu et al. (2021), Informer Zhou et al. (2021), TFT Lim et al. (2021), FEDformer Zhou et al. (2022), and linear models Zeng et al. (2023).
The experimental setup followed the guidelines outlined in Nie et al. (2022); Chen et al. (2023). We evaluated the models under long-term time series forecasting settings, with prediction lengths of 96, 192, 336, and 720. The evaluation metrics employed were mean square error (MSE) and mean absolute error (MAE). The dataset was divided into training, validation, and testing subsets, as suggested in Nie et al. (2022); Chen et al. (2023); Zeng et al. (2023). The input length was set to 512 as per Chen et al. (2023), and other hyper-parameters were determined based on the validation set results. The settings for the hyper-parameters are provided with the source code, which can be downloaded from The Author(2023) (s). We employed the AdamW optimizer Loshchilov and Hutter (2017) with the objective of minimizing the mean square error. The results are presented in Table 3.
Table 3: Performance comparisons. We bold the best-performing results.
@@ A.6 Long-Term Time Series Forecasting
The results suggest that our proposed method performs comparably to both TSMixer and PatchTST, as confirmed by a $t$-test with $\alpha=0.05$. This implies that the proposed method also achieves state-of-the-art performance on the long-term time series forecasting tasks. Chen et al. Chen et al. (2023) noted that the cross-variate information might not be as significant in these seven datasets. Therefore, the fact that our proposed method achieves state-of-the-art performance indicates that its components for capturing cross-variate information do not degrade performance, even when the cross-variate information is not relevant for the task. This suggests that our method is robust against irrelevant cross-variate information.

### v4

@@ A.5. Ablation Study
We have produced correlation-error diagrams employing all three measurements (MAE, RMSE, and MAPE). In Section 5.4, we showcase the correlation-error diagram with MAE in Fig. 10. This section introduces the remaining two correlation-error diagrams, namely, the correlation-error diagram with RMSE in Fig. 13, and the correlation-error diagram with MAPE in Fig. 14. The conclusions remain consistent. The random projection layer aids the proposed method in achieving greater diversity.
Figure 13. The correlation-error diagram illustrates the interplay of diversity and error (RMSE) across individual blocks in the model.
Figure 14. The correlation-error diagram illustrates the interplay of diversity and error (MAPE) across individual blocks in the model.
@@
Figure 15. The ablation study result measured using per-step MAE, RMSE, and MAPE. The random projection layer benefits the prediction for later time steps more compared to earlier time steps.
A.6. Runtime and Space Complexity
The proposed RPMixer’s efficiency closely aligns with that of the TSMixer, as both employ a similar all-MLP structure. With this in mind, we have conducted experiments to compare the runtimes of both the TSMixer and the proposed method. Table 3 shows the average runtime for each epoch.
Table 3. Runtime in seconds.
Data SD GBA GLA CA
TSMixer 196 591 954 2115
RPMixer 72 179 261 589
RPMixer w/o Fourier 46 85 106 231
RPMixer w/o random 69 168 250 568
The TSMixer is slower comparing to RPMixer due to its incorporation of batch normalization layers. When examining the different variants of RPMixer, we observe that the complex linear layer has a longer runtime compared to the random projection layer. Regarding the space complexity of the intermediate representation for all methods enumerated in Table 3, it is $O(nm)$, where $n$ denotes the number of nodes, and $m$ denotes the length of the input time series.
A.7. Long-Term Time Series Forecasting
The effectiveness of the proposed method in spatial-temporal forecasting tasks has been demonstrated. However, as a general time series forecasting method, we sought to assess its performance against other forecasting methods. To this end, we evaluated our method on seven multivariate long-term time series forecasting datasets (i.e., ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity, and Traffic), and compared it with two state-of-the-art methods, TSMixer (Chen et al., 2023) and PatchTST (Nie et al., 2022), which have shown superior performance over alternatives such as Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021), TFT (Lim et al., 2021), FEDformer (Zhou et al., 2022), and linear models (Zeng et al., 2023).
The experimental setup followed the guidelines outlined in (Nie et al., 2022; Chen et al., 2023). We evaluated the models under long-term time series forecasting settings, with prediction lengths of 96, 192, 336, and 720. The evaluation metrics employed were mean square error (MSE) and mean absolute error (MAE). The dataset was divided into training, validation, and testing subsets, as suggested in (Nie et al., 2022; Chen et al., 2023; Zeng et al., 2023). The input length was set to 512 as per (Chen et al., 2023), and other hyper-parameters were determined based on the validation set results. The settings for the hyper-parameters are provided with the source code, which can be downloaded from (The Author(2023), s). We employed the AdamW optimizer (Loshchilov and Hutter, 2017) with the objective of minimizing the mean square error. The results are presented in Table 4. The performance for TSMixer and PatchTST were obtained from their respective papers (Nie et al., 2022; Chen et al., 2023).
Table 4. Performance comparisons. We bold the best-performing results.
@@ A.7. Long-Term Time Series Forecasting
The results suggest that our proposed method performs comparably to both TSMixer and PatchTST, as confirmed by a $t$-test with $\alpha=0.05$. This implies that the proposed method also attains state-of-the-art performance on the long-term time series forecasting tasks. Chen et al. (Chen et al., 2023) noted that the cross-variate information might not be as significant in these seven datasets. However, it has been demonstrated in (Liu et al., 2023a) that effectively capturing the cross-variate information using an attention mechanism can enhance performance. This observation suggests that merging the random projection concept with an attention mechanism could be a promising future direction. It should be noted that the dimension count of most long-term time series forecasting datasets is significantly lower than that of large-scale traffic datasets. The modest performance of the proposed method for these datasets aligns with the results presented in Table 1. The dataset with the fewest dimensions exhibits the least improvement compared to the baseline methods.
A.8. Comparative Analysis of RPMixer and Alternative Forecasting Methods
The experimental section primarily compared our proposed RPMixer method with baseline methods that have reported results on the LargeST dataset (Liu et al., 2023b) available at the time of writing. TSMixer (Chen et al., 2023) is the only deep learning method that deviates from this rule, as it significantly inspired the RPMixer method. This section offers an analysis of alternative large-scale spatial-temporal forecasting methods and their potential integration with RPMixer.
A.8.1. Spatial-Temporal Forecasting Model
The PDFormer (Jiang et al., 2023a) model, a spatial-temporal graph neural network, employs a stack of transformer encoders with skip connections to process spatial-temporal data. Each encoder comprises three attention modules designed to capture long-range spatial dependencies, short-range spatial dependencies, and temporal information within a node. The random projection concept from our paper could be integrated into PDFormer by adding a fourth module, the spatial mixer, to the transformer encoder to diversify the intermediate representation output. MegaCRN (Jiang et al., 2023b), based on a graph convolutional recurrent unit (GCRU) encoder-decoder architecture, uses a meta-graph learner to generate a meta-graph for the recurrent unit. As MegaCRN is GCRU-based, it would be challenging to adopt a random projection layer to enhance its performance. Spatial and temporal identities are learnable embedding features (Shao et al., 2022b) that could be combined with different model architectures. Thus, it would be interesting to explore how they could further enhance RPMixer’s performance in large-scale spatial-temporal forecasting.
A.8.2. General Forecasting Model
Two significant components proposed in (Li et al., 2023b), reversible normalization (RevIN) and channel independence (CI), aim to address linear layers’ limitations. RevIN assists linear models in trend prediction, and CI aids linear models in handling multivariate time series with varying periods across different dimensions (or channels). RevIN could be incorporated into RPMixer to boost its trend prediction capability, while CI may not be compatible with RPMixer, which assumes dimensional dependency. The choice between CI and RPMixer should be based on the dataset’s characteristics. Factorized multilayer perceptrons (MLPs), proposed in (Li et al., 2023c) and evaluated on common multivariate time series forecasting benchmark datasets, model the dimension and temporal interaction of multivariate time series. It would be interesting to evaluate RPMixer’s performance after substituting the temporal mixer with a factorized MLP on large-scale spatial-temporal datasets.
The frequency-domain MLP (Yi et al., 2024) is a model that processes input multivariate time series with MLP in the frequency domain, using a frequency channel learner and a frequency temporal learner, and does not incorporate mixer blocks. The frequency temporal learner could be used as the alternative design for the temporal mixer used in RPMixer. While the FFT design is not our primary contribution, it offers an interesting perspective to assess if replacing FFT with decomposition (Zhou et al., 2022) or downsampling (Zhang et al., 2022) could enhance the handling of periodic data. In future work, we could weave these methods into RPMixer in numerous ways, either incorporating them into every temporal mixer or into the network’s input or output.
The N-BEATS (Oreshkin et al., 2019) model is a univariate time series forecasting model composed of a stack of residual building blocks, each a MLP with a forecast output and a backcast output. As a univariate time series model, its design principles could be used to refine the temporal model design in RPMixer.
A.9. Additional Benchmark Result
In this section, we present additional experimental results comparing the proposed RPMixer with another MLP model, Spatial and Temporal IDentity information (STID) (Shao et al., 2022b), on LargeST datasets (Liu et al., 2023b). We predominantly adopted the hyper-parameter settings from (Shao et al., 2022b, c). The experimental results, averaged over the full 12 time horizon, are displayed in Table 5.
Table 5. Comparisons between RPMixer and STID.
Data Method Param MAE RMSE MAPE
SD RPMixer 1.5M 16.90 27.97 11.07
STID 127K 29.05 47.73 19.28
GBA RPMixer 2.3M 19.06 31.54 15.09
STID 180K 31.20 48.72 26.11
GLA RPMixer 3.2M 18.46 30.13 11.34
STID 227K 31.86 50.51 20.20
CA RPMixer 7.8M 17.50 28.90 13.03
STID 380K 29.18 46.78 21.40
Overall, RPMixer surpasses STID in all error measurements. In terms of runtime, STID’s duration is roughly half that of RPMixer, due to STID’s smaller model size. Nevertheless, STID’s key innovation, i.e., learnable spatial and temporal embeddings (Shao et al., 2022b), is versatile and can be integrated with RPMixer. We performed an ablation study to assess the impact of the STID components. The results are available in Table 6.
Table 6. Ablation study of RPMixer with STID component.
Data Method Param MAE RMSE MAPE
SD RPMixer 1.5M 16.90 27.97 11.07
+ STID 2.6M 16.91 27.72 11.04
+ STID - random - Fourier 1.6M 17.71 28.64 11.98
GBA RPMixer 2.3M 19.06 31.54 15.09
+ STID 3.6M 19.07 31.02 15.74
+ STID - random - Fourier 3.3M 20.58 32.45 18.24
GLA RPMixer 3.2M 18.46 30.13 11.34
+ STID 4.7M 18.73 30.06 11.75
+ STID - random - Fourier 5.4M 20.04 31.62 13.44
CA RPMixer 7.8M 17.50 28.90 13.03
+ STID 9.7M 17.55 28.74 13.22
+ STID - random - Fourier 14.9M 18.91 30.05 16.09
In this ablation study, we considered three settings: 1) RPMixer, the method introduced in this paper, 2) RPMixer + STID, which is RPMixer enhanced with the learnable spatial and temporal embeddings from STID, and 3) RPMixer + STID - Random - Fourier, the previous setting with the main components (i.e., random projection layers and complex linear layers) of RPMixer turned off. The first setting serves as the baseline for this study, the second setting demonstrates the potential benefit of the STID components on the proposed method, and the third setting illustrates the relative importance of the STID components versus the RPMixer components. By integrating STID into RPMixer (i.e., RPMixer + STID), the RMSE on all datasets improves. However, when the RPMixer components are turned off (i.e., RPMixer + STID - Random - Fourier), the performance on all three measures deteriorates. The RPMixer components have a greater impact on performance compared to the STID components.
The number of parameters for RPMixer + STID is higher compared to RPMixer, due to the inclusion of additional embedding layers. In terms of runtime, incorporating STID increases the runtime by an average of 22% because of these extra layers. Overall, integrating STID components into RPMixer could be a promising strategy to further enhance the performance of RPMixer.
A.10. Interpretability and Ease of Use
Regarding interpretability, RPMixer does not compute adjacency matrices like many baseline methods, thus it does not provide intrinsic interpretability. However, post-hoc interpretability can be achieved through methods like sensitivity analysis (Wang et al., 2019; Ismail Fawaz et al., 2019; Wang et al., 2021; Yeh et al., 2023b), which can highlight the more important parts of the input. In terms of ease of use, RPMixer is composed of four common layers: linear, ReLU, FFT, and iFFT, which can be found in almost all deep learning libraries. Therefore, we believe RPMixer is relatively easy to implement.